OpenStudy (ilovecomputers1):

What exponential function is the best fit for the data in the table? x f(x) 1 -4 3 -1 4 3

OpenStudy (ilovecomputers1):

f(x) = 4(3)^x - 1 + 4 f(x) = 4(3^)x - 1 - 4 f(x) = 1/4(3)^x - 1 + 4 f(x) = 1/4 (3)^x - 1 - 4